Given that A=2i-3j+6k and B=i-2j+2k.find the scalar projection of B onto A


1
Expert's answer
2021-11-29T11:41:38-0500

The scalar projection of B onto A "=(\vec a\cdot\vec b)/|\vec a|="


"=(2\cdot1+(-3)\cdot(-2)+6\cdot2)/\sqrt{2^2+3^2+6^2}="


"=20\/7=2.86" . Answer